OpenStudy (anonymous):

Simplify 3 sqrt54 - 2 sqrt24 - sqrt 96 + 4 sqrt 63

OpenStudy (anonymous):

\[3\sqrt{54} - 2\sqrt{24} - \sqrt{96} + 4\sqrt{63}\]

OpenStudy (anonymous):

\[\sqrt{6}+12\sqrt{7}\]

OpenStudy (anonymous):

what are the steps ??

OpenStudy (anonymous):

\[3\sqrt{69*6}-2\sqrt{4*6}-\sqrt{16*6}+4\sqrt{9*7}\]

OpenStudy (anonymous):

\[3\sqrt{9*6}\]